# Doubly Robust Alignment for Large Language Models

### v1

@@ B DRPO Algorithm Details and Practical Implementation
$\mathcal{J}(\pi_{\theta};\widehat{\pi}_{\mathrm{ref}},\widehat{g}_{\eta},\mathcal{D})=\widehat{p}_{\textrm{DR}}(\pi)-\beta\mathbb{E}_{X\sim\mathcal{D}}[{D}_{\mathrm{KL}}[\pi(\cdot\mid X)\,\|\,\widehat{\pi}_{\mathrm{ref}}(\cdot\mid X)]].$ (37)
@@
$\nabla_{\theta}\mathcal{J}(\pi_{\theta})$ $=\frac{1}{2}\mathbb{E}_{X,Y^{(1)},Y^{(2)}\sim\mathcal{D}}\left[\sum_{a=1}^{2}\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid X)}\left[\widehat{g}(X,y,Y^{(a)})\nabla_{\theta}\log\pi_{\theta}(y|X)\right]\right.$
$\left.+\sum_{a=1}^{2}(-1)^{a-1}\frac{\nabla_{\theta}\pi_{\theta}(Y^{(a)}|X)}{\widehat{\pi}_{\mathrm{ref}}(Y^{(a)}|X)}\big{(}Z-\widehat{g}(X,Y^{(1)},Y^{(2)})\big{)}\right]$
$-\beta\nabla_{\theta}{D}_{\mathrm{KL}}[\pi_{\theta}(\cdot\mid X)\,\|\,\widehat{\pi}_{\mathrm{ref}}(\cdot\mid X)]$ (38)
@@
The empirical loss function is constructed such that its negative gradient corresponds to $\nabla_{\theta}\mathcal{J}(\pi_{\theta})$ in Equation B. The direct-method term is approximated using Monte Carlo sampling by drawing several new responses $\mathcal{D}_{X}^{*}:=\{Y^{*}\mid Y^{*}\sim\pi_{\theta}(\cdot\mid X)\}$ from the current policy $\pi_{\theta}$ for a given prompt X at each policy update. A k3-type empirical KL divergence is utilized, following [Shao et al., , 2024].
@@
$\mathcal{L}_{\textup{DRPO}}$ $=$ $-\frac{1}{2}\mathbb{E}_{X,Y^{(1)},Y^{(2)}\sim\widetilde{\mathcal{D}}}\Bigg{[}\underbrace{\mathbb{E}_{Y^{*}\sim\mathcal{D}_{X}^{*}}\left[\widehat{g}(Y^{*},Y^{(2)},X)\log\pi_{\theta}(Y^{*}|X)\right]}_{\mathrm{term\ I}}$ (40)
$+\mathrm{sg}\bigg{(}\underbrace{\mathrm{clip}\Big{(}\frac{\pi_{\theta}(Y^{(1)}|X)}{\pi_{\mathrm{ref}}(Y^{(1)}|X)},1-\epsilon_{1},1+\epsilon_{2}\Big{)}\big{(}Z-\widehat{g}(Y^{(1)},Y^{(2)},X)\big{)}}_{\mathrm{term\ II}}\bigg{)}\log\pi_{\theta}(Y^{(1)}\mid X)\Bigg{]}$
@@
where $\mathrm{sg}(\cdot)$ denotes stop-gradient operation, $\mathrm{clip}(\cdot,a,b)$ signifies clipping values to the interval $[a,b]$, and $\epsilon_{1},\epsilon_{2}$ are hyperparameters defining the clipping range. See Algorithm 1 for details.
@@
0:  reference policy $\widehat{\pi}_{\mathrm{ref}}$, preference model $\widehat{g}$, offline dataset $\widetilde{\mathcal{D}}=\{X_{i},Y^{(1)}_{i},Y^{(2)}_{i},Z_{i}\}$, clipping range $[\epsilon_{1},\epsilon_{2}]$, regularization parameter $\beta$, and other hyperparameters, effective batch size $|\mathcal{B}|$, learning rate $\alpha$ and the optimizer, number of Monte Carlo samples $|\mathcal{D}^{*}|$.
@@
1:  Initialize policy $\pi_{\theta}^{(0)}$, total train steps $T=\frac{|\widetilde{\mathcal{D}}|}{|\mathcal{B}|}$. For brevity let number of training epochs $N=1$.
@@
4:         Sample $\mathcal{D}^{*}_{X_{i}}=\{Y^{*}_{j}\mid Y^{*}_{j}\sim\pi^{(t-1)}_{\theta}(\cdot\mid X_{i})\}_{j\in[|\mathcal{D}^{*}|]}$.
5:         Esitmate term I: $\widehat{\mathrm{I}}_{i}=\frac{1}{|\mathcal{D}^{*}_{X_{i}}|}\sum_{Y^{*}\in\mathcal{D}^{*}_{X_{i}}}\widehat{g}(Y^{*},Y_{i}^{(2)},X_{i})\log\pi^{(t-1)}_{\theta}(Y^{*}|X_{i})$;
@@
$\widehat{\mathrm{II}}_{i}=\mathrm{clip}\Bigg{(}\frac{\pi^{(t-1)}_{\theta}(Y^{(1)}_{i}|X_{i})}{\pi_{\mathrm{ref}}(Y^{(1)}_{i}|X_{i})},1-\epsilon_{1},1+\epsilon_{2}\Bigg{)}\big{(}Z-\widehat{g}(Y^{(1)}_{i},Y^{(2)}_{i},X_{i})\big{)};$
@@
$\widehat{D}_{\mathrm{KL}_{i}}=\frac{1}{|\mathcal{D}^{*}_{X_{i}}|}\sum_{Y^{*}\in\mathcal{D}^{*}_{X_{i}}}\left(\frac{\widehat{\pi}_{\mathrm{ref}}(Y^{*}\mid X)}{\pi_{\theta}(Y^{*}\mid X)}-1-\log\frac{\widehat{\pi}_{\mathrm{ref}}(Y^{*}\mid X)}{\pi_{\theta}(Y^{*}\mid X)}\right);$
@@
10:     update $\theta^{(t)}$ with gradient descent and get $\pi_{\theta}^{(t)}$: $\theta^{(t)}=\theta^{(t-1)}-\alpha\nabla_{\theta}\mathcal{L}$;
@@ C Experiments Implementation details
For the baseline models training, we follow the framework of TRL: Transformer Reinforcement Learning [von Werra et al., , 2020] and Transformers: State-of-the-Art Natural Language Processing [Wolf et al., , 2020]. For the general preference model, we follow the framework of general-preference/general-preference-model proposed by Zhang et al., 2024d . All models were trained with default hyperparameter configurations unless otherwise specified.
The Preference Evaluation experiments are conducted on a machine equipped with an NVIDIA RTX 6000 Ada GPU and an AMD Ryzen Threadripper PRO 7945WX 12-core CPU. The Preference Optimization experiments are performed on a system with an H20 NVLink GPU and a 20 vCPU Intel(R) Xeon(R) Platinum 8457C processor. AdamW [Loshchilov and Hutter, , 2019] are used as default optimizer.
@@ C.1 Preference Evaluation Experiment on IMDb
Oracle Preference Model. Since the IMDb dataset does not contain human preference labels, we adopt the known sentiment classifier siebert/sentiment-roberta-large-english [Hartmann et al., , 2023], as a ground-truth reward-based labeler. This classifier will give a score $s(X,Y)=p(\text{positive}\ |\ X,Y)$, which we convert into a reward signal using the log-odds transformation:
@@
where $\sigma(\cdot)$ is the sigmoid function.
@@
Data Generation and Policy Training Process. We begin by fine-tuning a supervised fine-tuning (SFT) model based on the EleutherAI/gpt-neo-125m base model [Black et al., , 2021] for 3 epochs using the 25,000 training samples from the IMDb dataset. Prompts are constructed by extracting 5-word prefixes from movie reviews. Using the fine-tuned SFT model as the reference policy, we generate pairs of completions for each prompt. Next, we use the oracle preference model to estimate the preference probabilities between each pair of completions. Based on these probabilities, we sample binary preference labels indicating which response is preferred. This synthetic preference dataset is then used to train a target policy using the Direct Preference Optimization (DPO) algorithm over an additional 3 epochs. To quantify the relative preference for the target policy over the reference policy, we adopt a Monte Carlo estimation approach. Specifically, for each of the 25,000 prefixes in the IMDb test set, both the target and reference policies generate a single completion. The oracle preference model is then used to compute the preference probability between the two completions. Aggregating these results, we estimate the overall probability, which is 0.681, that the target policy’s outputs are preferred over those of the reference policy.
@@
Preference Evaluation Process. We consider two versions of the reference policy estimator $\widehat{\pi}_{\textrm{ref}}$: a correctly specified version, where $\widehat{\pi}_{\textrm{ref}}$ corresponds to the SFT model, and a misspecified version, where $\widehat{\pi}_{\textrm{ref}}$ corresponds to the untrained base model. Similarly, we consider two versions of the preference estimator $\widehat{g}$: a correctly specified version, which uses the oracle preference model, and a misspecified version, where $\widehat{g}$ is drawn uniformly at random from $[0,1]$. By taking all pairwise combinations of $\widehat{\pi}_{\textrm{ref}}$ and $\widehat{g}$, we construct four distinct variants of the preference evaluation framework. For the Direct Method (DM) estimator in Equation 6, we apply a Monte Carlo approach by sampling 8 responses from the target policy for each prompt. For the Importance Sampling (IS) estimator in Equation 7, we use a clipping ratio of 100 when $\widehat{\pi}_{\textrm{ref}}$ is correctly specified and 40 when it is misspecified. In contrast to the clipping ratio used during preference optimization, a larger ratio is adopted here to better demonstrate the double robustness property of our preference evaluation framework.
C.2 Preference Optimization Experimen on Real Data
Baseline models training. For the summarization task, we adopt models from a group of Hugging Face, cleanrl, known for their validated and quality-assured implementations [Huang et al., , 2024]. Specifically, we use cleanrl/EleutherAI_pythia-1b-deduped__sft__tldr as both the reference and initial policy model. This SFT policy is trained via token-level supervised fine-tuning on human-written summaries from a filtered TL;DR Reddit dataset [Huang et al., , 2024]. The associated reward model is
@@ C.2 Preference Optimization Experimen on Real Data
cleanrl/EleutherAI_pythia-1b-deduped__reward__tldr. For Proximal Policy Optimization (PPO) training, we search the hyperparameter over the KL coefficient $\beta\in\{0.05,0.1,0.2\}$ and select $\beta=0.05$ based on empirical performance. Notably,we observe that PPO training can experience policy collapse under low-precision, as the value function fails to fit accurately; thus, PPO models are trained under full precision (FP32). In contrast, all our models are trained using bfloat16 (BF16) for improved computational efficiency. To ensure a fair comparison, we set the maximum response length to 128 for all models, providing a consistent basis for assessing summarization quality.
[…]
DRPO Implementation. DRPO implementation inherits transformers.Trainer class. For DRPO-BT, we compute the rewards for two candidate responses and output the preference probability under the BT framework as $\widehat{g}$. For DRPO-GPM, we directly compute the preference probability using the corresponding general preference model [Zhang et al., 2024d, ]. Although our proposed algorithm allows the use of a more powerful general preference model for estimating $\widehat{g}$, as in [Munos et al., , 2023], we ensure fairness by training all preference models using the same base model and dataset. This avoids introducing any additional information that could bias the comparison. For both tasks, we set the clipping range to $[0.04,2.5]$, a fairly casual (and wide) specification only to force the IS ratio to not deviate far from 1 and thus not inject too much variance into our estimation. The regularization parameter $\beta$ is set to 0.04, the same as that in the default trl implementation for GRPO [Shao et al., , 2024], which also uses k3-type empirical KL divergence. The number of Monte Carlo samples $|\mathcal{D}^{*}|$ is set to 3 (TL;DR) or 2 (HH). Although more samples may mitigate bias, the effect of adding samples is marginally decreasing (since the convergence rate is $O((n^{*})^{-\frac{1}{2}})$). As such, it is proper to choose a parsimonious volume of samples and thus incurring little extra computational cost compared to PPO. Other not-mentioned hyperparameters are simply set to default values. For further details, please refer to the examples in the codebase.
Evaluation. We compare DRPO with DPO and PPO using GPT-4o-mini to evaluate the quality of generated response of each task. Specifically, for the language model fine-tuned by either baseline or our method, we can sample a response at a certain temperature after it receives a prompt. With the responses of two methods (say A and B), we feed them with a query asking GPT to judge which is more aligned with certain demands. The query template used for TL;DR is shown in Table 2, which tries to avoid GPT’s favor of lengthy responses following [Ye et al., , 2025]. The query template used for HH is shown in Table 3, a standard template that is widely adopted by e.g. [Rafailov et al., , 2023, Wu et al., 2024a, , Ye et al., , 2025]. It is noteworthy that we randomly shuffle the order of the responses for each query to eliminate the potential bias from the order of the responses.
@@
D Additional Empirical Results
In this section, we first provide pairwise win rates on the TL;DR dataset with other sampling temperatures (see Figure 5). Our method consistently dominates across all temperatures. DPO’s performance improves when temperature gets lower, which is in line with results in [Rafailov et al., , 2023]. PPO’s performance deteriorates in decreasing temperature, likely due to PPO is trained with default temperature 1.0. Next, we present pairwise win rates on HH dataset with other sampling temperatures (see Figure 6). The results are consistent with that of temperature 1.0. In general, DRPO-GPM $\succ$ DRPO-BT $\approx$ DPO $\succ$ PPO, showcasing the robustness of our algorithm.
Additionally, we present some of the sampled responses of our method and baselines and how gpt-4o-mini judges the quality of the completions. See Table 4, 5, 6, 7 for TL;DR examples and Table 8, 9, 10, 11 for HH examples.

### v2

@@ B DRPO Algorithm Details and Practical Implementation
$\mathcal{J}(\pi_{\theta};\widehat{\pi}_{\mathrm{ref}},\widehat{g}_{\eta},\mathcal{D})=\widehat{p}_{\textrm{DR}}(\pi)-\beta\mathbb{E}_{X\sim\mathcal{D}}{D}_{\mathrm{KL}}[\pi(\bullet\mid X)\,\|\,\widehat{\pi}_{\mathrm{ref}}(\bullet\mid X)].$ (38)
@@
$\nabla_{\theta}\mathcal{J}(\pi_{\theta})$ $=$ $\frac{1}{2}\mathbb{E}_{X,Y^{(1)},Y^{(2)}\sim\mathcal{D}}\left\{\sum_{a=1}^{2}\mathbb{E}_{y\sim\pi_{\theta}(\bullet\mid X)}\left[\widehat{g}(X,y,Y^{(a)})\nabla_{\theta}\log\pi_{\theta}(y|X)\right]\right$. (39)
$\left.+\sum_{a=1}^{2}(-1)^{a-1}\frac{\nabla_{\theta}\pi_{\theta}(Y^{(a)}|X)}{\widehat{\pi}_{\mathrm{ref}}(Y^{(a)}|X)}\big(Z-\widehat{g}(X,Y^{(1)},Y^{(2)})\big)\right\}$
$-\beta\nabla_{\theta}{D}_{\mathrm{KL}}[\pi_{\theta}(\bullet\mid X)\,\|\,\widehat{\pi}_{\mathrm{ref}}(\bullet\mid X)]$
@@
The empirical loss function is constructed such that its negative gradient corresponds to $\nabla_{\theta}\mathcal{J}(\pi_{\theta})$ in Equation 39. The direct-method term is approximated using Monte Carlo sampling by drawing several new responses $\mathcal{D}_{X}^{*}:=\{Y^{*}\mid Y^{*}\sim\pi_{\theta}(\bullet\mid X)\}$ from the current policy $\pi_{\theta}$ for a given prompt X at each policy update. A k3-type empirical KL divergence is utilized, following [43].
@@
$\mathcal{L}_{\textup{DRPO}}$ $=$ $-\frac{1}{2}\mathbb{E}_{X,Y^{(1)},Y^{(2)}\sim\widetilde{\mathcal{D}}}\Bigg\{\underbrace{\mathbb{E}_{Y^{*}\sim\mathcal{D}_{X}^{*}}\left[\widehat{g}(Y^{*},Y^{(2)},X)\log\pi_{\theta}(Y^{*}|X)\right]}_{\mathrm{term\ I}}$ (41)
$+\mathrm{sg}\bigg(\underbrace{\mathrm{clip}\Big(\frac{\pi_{\theta}(Y^{(1)}|X)}{\pi_{\mathrm{ref}}(Y^{(1)}|X)},1-\epsilon_{1},1+\epsilon_{2}\Big)\big(Z-\widehat{g}(Y^{(1)},Y^{(2)},X)\big)}_{\mathrm{term\ II}}\bigg)\log\pi_{\theta}(Y^{(1)}\mid X)\Bigg\}$
@@
where $\mathrm{sg}(\bullet)$ denotes stop-gradient operation, $\mathrm{clip}(\bullet,a,b)$ signifies clipping values to the interval $[a,b]$, and $\epsilon_{1},\epsilon_{2}$ are hyperparameters defining the clipping range. See full details in Algorithm 1
@@
0: reference policy $\widehat{\pi}_{\mathrm{ref}}$, preference model $\widehat{g}$, offline dataset $\widetilde{\mathcal{D}}=\{X_{i},Y^{(1)}_{i},Y^{(2)}_{i},Z_{i}\}$, clipping range $\epsilon_{1},\epsilon_{2}]$, regularization parameter $\beta$, and other hyperparameters, effective batch size $|\mathcal{B}|$, learning rate $\alpha$ and the optimizer, number of Monte Carlo samples $|\mathcal{D}^{*}|$.
@@
1:Initialize policy $\pi_{\theta}^{(0)}$, total train steps $T=\frac{|\widetilde{\mathcal{D}}|}{|\mathcal{B}|}$. For brevity let the number of training epochs $N=1$.
@@
4:   Sample $\mathcal{D}^{*}_{X_{i}}=\{Y^{*}_{j}\mid Y^{*}_{j}\sim\pi^{(t-1)}_{\theta}(\bullet\mid X_{i})\}_{j\in[|\mathcal{D}^{*}|]}$.
5:   Esitmate term I:
$\widehat{\mathrm{I}}_{i}=\frac{1}{|\mathcal{D}^{*}_{X_{i}}|}\sum_{Y^{*}\in\mathcal{D}^{*}_{X_{i}}}\widehat{g}(Y^{*},Y_{i}^{(2)},X_{i})\log\pi^{(t-1)}_{\theta}(Y^{*}|X_{i})$
@@
$\widehat{\mathrm{II}}_{i}=\mathrm{clip}\Bigg(\frac{\pi^{(t-1)}_{\theta}(Y^{(1)}_{i}|X_{i})}{\pi_{\mathrm{ref}}(Y^{(1)}_{i}|X_{i})},1-\epsilon_{1},1+\epsilon_{2}\Bigg)\big(Z-\widehat{g}(Y^{(1)}_{i},Y^{(2)}_{i},X_{i})\big)$
@@
$\widehat{D}_{\mathrm{KL}_{i}}=\frac{1}{|\mathcal{D}^{*}_{X_{i}}|}\sum_{Y^{*}\in\mathcal{D}^{*}_{X_{i}}}\left(\frac{\widehat{\pi}_{\mathrm{ref}}(Y^{*}\mid X)}{\pi_{\theta}(Y^{*}\mid X)}-1-\log\frac{\widehat{\pi}_{\mathrm{ref}}(Y^{*}\mid X)}{\pi_{\theta}(Y^{*}\mid X)}\right)$
@@
10:  update $\theta^{(t)}$ with gradient descent and get $\pi_{\theta}^{(t)}$:
$\theta^{(t)}=\theta^{(t-1)}-\alpha\nabla_{\theta}\mathcal{L}$
@@ C Experiments Implementation details
For the baseline models training, we follow the framework of TRL: Transformer Reinforcement Learning [155] and Transformers: State-of-the-Art Natural Language Processing [163]. For the general preference model, we follow the framework of general-preference/general-preference-model proposed by Zhang et al. [19]. All models were trained with default hyperparameter configurations unless otherwise specified.
The Preference Evaluation experiments are conducted on a machine equipped with an NVIDIA RTX 6000 Ada GPU and an AMD Ryzen Threadripper PRO 7945WX 12-core CPU. The Preference Optimization experiments are performed on a system with an H20 NVLink GPU and a 20 vCPU Intel(R) Xeon(R) Platinum 8457C processor. AdamW [164] are used as default optimizer.
@@ C.1 Preference Evaluation Experiment on IMDb
Oracle Preference Model. Since the IMDb dataset does not contain human preference labels, we adopt the known sentiment classifier siebert/sentiment-roberta-large-english [165], as a ground-truth reward-based labeler. This classifier will give a score $s(X,Y)=p(\text{positive}\ |\ X,Y)$, which we convert into a reward signal using the log-odds transformation:
@@
where $\sigma(\bullet)$ is the sigmoid function.
@@
Data Generation and Policy Training Process. We begin by fine-tuning supervised fine-tuning (SFT) models initialized from two base architectures of different scales: the EleutherAI/gpt-neo-125m [154] and the Qwen/Qwen2.5-7B [166]. Both models are trained for three epochs on 25,000 samples from the IMDb training set. Prompts are constructed by extracting 5-word prefixes from movie reviews. Using the fine-tuned SFT model as the reference policy, we generate pairs of completions for each prompt. Next, we use the oracle preference model to estimate the preference probabilities between each pair of completions. Based on these probabilities, we sample binary preference labels indicating which response is preferred. This synthetic preference dataset is then used to train a target policy using the Direct Preference Optimization (DPO) algorithm over an additional 3 epochs. To quantify the relative preference for the target policy over the reference policy, we adopt a Monte Carlo estimation approach. Specifically, for each of the 25,000 prefixes in the IMDb test set, both the target and reference policies generate a single completion. The oracle preference model is then used to compute the preference probability between the two completions. Aggregating these results, we estimate the overall probability, which is 0.681, that the target policy’s outputs are preferred over those of the reference policy.
@@
Preference Evaluation Process. We consider two versions of the reference policy estimator $\widehat{\pi}_{\textrm{ref}}$: a correctly specified version, where $\widehat{\pi}_{\textrm{ref}}$ corresponds to the SFT model, and a misspecified version, where $\widehat{\pi}_{\textrm{ref}}$ corresponds to the untrained base model. Similarly, we consider two versions of the preference estimator $\widehat{g}$: a correctly specified version, which uses the oracle preference model, and a misspecified version, where $\widehat{g}$ is drawn uniformly at random from $[0,1]$. By taking all pairwise combinations of $\widehat{\pi}_{\textrm{ref}}$ and $\widehat{g}$, we construct four distinct variants of the preference evaluation framework. For the Direct Method (DM) estimator in Equation 6, we apply a Monte Carlo approach by sampling 8 responses from the target policy for each prompt. For the Importance Sampling (IS) estimator in Equation 7, we use a clipping ratio of 100 when $\widehat{\pi}_{\textrm{ref}}$ is correctly specified and 40 when it is misspecified. In contrast to the clipping ratio used during preference optimization, a larger ratio is adopted here to better demonstrate the double robustness property of our preference evaluation framework. The results based on the EleutherAI/gpt-neo-125m model [154] are presented in Figure 3 in Section 6, while those based on the Qwen/Qwen2.5-7B model [166] are summarized in Table 4.
Table 4: MSE of the proposed preference estimator with a 7B base model. The preference model and reference policy can be misspecified or correctly specified.
Sample size 500 1000 2000 3000
Both correct 0.002212 0.001160 0.000702 0.000390
Wrong preference model 0.024942 0.018757 0.016763 0.016594
Wrong reference model 0.066897 0.021389 0.013358 0.008383
Both wrong 0.265155 0.069340 0.043276 0.045954
C.2 Preference Optimization Experiment on Real Data
@@
Baseline models training. For the summarization task, we adopt models from a group of Hugging Face, cleanrl, known for their validated and quality-assured implementations [153]. Specifically, we use cleanrl/EleutherAI_pythia-1b-deduped__sft__tldr as both the reference and initial policy model. This SFT policy is trained via token-level supervised fine-tuning on human-written summaries from a filtered TL;DR Reddit dataset [153]. The associated reward model is cleanrl/EleutherAI_pythia-1b-deduped__reward__tldr. For PPO training, we search the hyperparameter over the KL coefficient $\beta\in\{0.05,0.1,0.2\}$ and select $\beta=0.05$ based on empirical performance. Notably,we observe that PPO training can experience policy collapse under low-precision, as the value function fails to fit accurately; thus, PPO models are trained under full precision (FP32). In contrast, all our models are trained using bfloat16 (BF16) for improved computational efficiency. To ensure a fair comparison, we set the maximum response length to 128 for all models, providing a consistent basis for assessing summarization quality. For DPO and its variants, we use default hyperparameter setting in TRL with BF16 precision. Notably, Dr.DPO had no official TRL implementation, so we adapt the loss function in DPOTrainer with Dr.DPO’s reweighting strategy and use the suggested hyperparameters in [54].
For human dialogue, the SFT model is trained from the base model Qwen/Qwen2.5-1.5B [166] to better align with the Helpfulness and Harmlessness (HH) dataset. Unlike the summarization SFT model, this version leverages both the preferred (chosen) and non-preferred (rejected) responses from the HH preference dataset. It is trained for 3 epochs. We also train three versions of the reward model, all from the same base model (Qwen/Qwen2.5-1.5B) to avoid additional information, corresponding to epochs 1, 2, and 3, as we observe that PPO training in this setting is highly sensitive to the reward model. When the reward model overfits or becomes overly confident, the KL penalty becomes ineffective, and PPO tends to suffer from policy collapse, hacking the reward model by repeating high-reward tokens. To mitigate this issue, we select the reward model from epoch 1, which achieves an evaluation accuracy of 72.1%. We further conduct a hyperparameter search over KL coefficients $\beta\in\{0.05,0.1,0.2\}$ and learning rates in $\{1\text{e-}7,1\text{e-}6,3\text{e-}6\}$. We select a KL coefficient of 0.05 combined with a learning rate of $1\text{e-}7$ as it yields the most stable and effective PPO training performance. Similar to those in summarization, DPO and its variants are trained with default setting.
@@
DRPO Implementation DRPO implementation inherits transformers.Trainer class. For DRPO-BT, we compute the rewards for two candidate responses and output the preference probability under the BT framework as $\widehat{g}$. For DRPO-GPM, we directly compute the preference probability using the corresponding general preference model [19]. Although our proposed algorithm allows the use of a more powerful general preference model for estimating $\widehat{g}$, as in [57], we ensure fairness by training all preference models using the same base model and dataset. This avoids introducing any additional information that could bias the comparison. For both tasks, we set the clipping range to $[0.04,2.5]$, a fairly casual (and wide) specification only to force the IS ratio to not deviate far from 1 and thus not inject too much variance into our estimation. The regularization parameter $\beta$ is set to 0.04, the same as that in the default trl implementation for GRPO [43], which also uses k3-type empirical KL divergence. The number of Monte Carlo samples $|\mathcal{D}^{*}|$ is set to 3 (TL;DR) or 2 (HH). Although more samples may mitigate bias, the effect of adding samples is marginally decreasing (since the convergence rate is $O((n^{*})^{-\frac{1}{2}})$). As such, it is proper to choose a parsimonious volume of samples and thus incurring little extra computational cost compared to PPO. Other not-mentioned hyperparameters are simply set to default values. For further details, please refer to the examples in the codebase.
Evaluation For in-distribution evaluation, we compare DRPO with DPO and PPO using GPT-4o-mini to evaluate the quality of generated response of each task. Specifically, for the language model fine-tuned by either baseline or our method, we can sample a response at a certain temperature after it receives a prompt. With the responses of two methods (say A and B), we feed them with a query asking GPT to judge which is more aligned with certain demands. The query template used for TL;DR is shown in Table 5, which tries to avoid GPT’s favor of lengthy responses following [40]. The query template used for HH is shown in Table 6, a standard template that is widely adopted by e.g. [26, 41, 40]. It is noteworthy that we randomly shuffle the order of the responses for each query to eliminate the potential bias from the order of the responses.
For out-of-distribution evaluation in HH dataset, we evaluate our models using the AlpacaEval 2.0 benchmark [159], an LLM-based automatic evaluator designed to assess models’ general performance. The prompt set in AlpacaEval 2.0 is derived from AlpacaFarm [167], which contains a broad collection of human-written instructions covering a wide range of general-purpose tasks beyond the Helpful–Harmless (HH) domain. By default, AlpacaEval 2.0 compares each model-generated response against a reference response produced by GPT-4-Turbo, and a GPT-4-Turbo-based annotator determines which of the two is preferred. However, we observed that all fine-tuning algorithms achieved consistently low win rates when evaluated against GPT-4-Turbo references, likely due to the substantial capability gap between GPT-4-Turbo and the fine-tuned models. To ensure a fairer and more interpretable comparison, we therefore replace the reference responses with those generated by the SFT model, allowing AlpacaEval 2.0 to compute the win rate of each fine-tuning algorithm relative to the SFT baseline.
@@
D Additional Empirical Results
In this section, we first provide pairwise win rates on the TL;DR dataset with other sampling temperatures (see Figure 5). Our method consistently dominates across all temperatures. DPO’s performance improves when temperature gets lower, which is in line with results in [26]. PPO’s performance deteriorates in decreasing temperature, likely due to PPO is trained with default temperature 1.0. Next, we present pairwise win rates on HH dataset with other sampling temperatures (see Figure 6). The results are consistent with that of temperature 1.0. In general, DRPO-GPM $\succ$ DRPO-BT $\approx$ DPO $\succ$ PPO, showcasing the robustness of our algorithm.
Additionally, we present some of the sampled responses of our method and baselines and how gpt-4o-mini judges the quality of the completions. See Table 7, 8, 9, 10 for TL;DR examples and Table 11, 12, 13, 14 for HH examples.